OpenStudy (anonymous):

Identify the equation of the circle that passes through (-3, -5) with center (4, -7). a.) (x - 4)2 + (y + 7)2 = sqrt 53 b.) (x + 4)2 + (y - 7)2 = sqrt 53 c.) (x + 4)2 + (y - 7)2 = 53 d.) (x - 4)2 + (y + 7)2 = 53

jimthompson5910 (jim_thompson5910):

what is the distance between the two given points?

OpenStudy (anonymous):

I graphed it and came up with 7 as the distance? Most likely it could be wrong.

OpenStudy (anonymous):

=(

jimthompson5910 (jim_thompson5910):

unfortunately it is incorrect

jimthompson5910 (jim_thompson5910):

you would use the distance formula to find the distance between the two points are you familiar with that formula?

OpenStudy (anonymous):

Yupp!

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

so plug each point into the formula and show me what you get

OpenStudy (anonymous):

I got the distance as the sqrt of 53 = 7.3 Is that okay?

jimthompson5910 (jim_thompson5910):

perfect, that's the distance

jimthompson5910 (jim_thompson5910):

so the radius is exactly \(\large r = \sqrt{53}\) units long

jimthompson5910 (jim_thompson5910):

the center in general is (h,k) so if the center is given to be (4, -7), then (h,k) = (4, -7) ---> h = 4, k = -7 we now plug this all into \(\large (x-h)^2+(y-k)^2=r^2\) to get \[\large (x-h)^2+(y-k)^2=r^2\] \[\large (x-4)^2+(y-(-7))^2=(\sqrt{53})^2\] \[\large (x-4)^2+(y+7)^2=53\] which is the equation of the circle

OpenStudy (anonymous):

Wow! Thank you so much! You are always very helpful! =)

jimthompson5910 (jim_thompson5910):

I'm glad I am, yw

OpenStudy (dan815):